OpenStudy (anonymous):

Find the area of the triangle when base = 15cm and height = 18cm. A. 125 cm^2 B. 129 cm^2 C. 133 cm^2 D. 135 cm^2

OpenStudy (anonymous):

\[\frac{1}{2}\times 15\times 18\]